OpenStudy (anonymous):

Which of the following is an equation of a line passing through the points (-6,1) and (3,-2) A. -3y=x+3 B. 3y=x-3 C. y-1=-1/3x D. y+1=1/3x

OpenStudy (the_fizicx99):

Find the slope, M = Y2 - Y1/X2-x1 Then plug into y-y1=m(x-x1)

OpenStudy (anonymous):

I got 3/9 for the slope, do i divide by 3? and I need help plugging in

OpenStudy (the_fizicx99):

M = -2-1/3-(-6) = -3/9 = -1/3 The slope is -1/3 y-y1=m(x-x1) y-1 = -1/3(x-(-6))

OpenStudy (anonymous):

ok, thx